Help Sev ob (5 points) $20,000 and a 25 percent tax on income between $20,000 and $30,000 Any income above $30000 is tased at 35 percemt Instructions: Round your answers to the nearest whole number a If your income is $70,000, how much will you pay in taxes? on the first S10,000 of income, but a 15 percent tax on earnings between $10,000 and b. Determine your marginal tax rate c Determine your average tax rate 1 of 1 8 9 5 6
Answer #1

a.

Amount of total tax for the income $70,000:

First $10,000 at 5% = $10,000 × 5% = $500

Income more than $10,000 to $20,000 at 15% = (20,000 – 10,000) × 15% = 10,000 × 15% = $1,500

Income more than $20,000 to $30,000 at 25% = (30,000 – 20,000) × 25% = 10,000 × 25% = $2,500

Income above $30,000 at 35% = (70,000 – 30,000) × 35% = 40,000 × 35% = $14,000

Total tax = 500 + 1,500 + 2,500 + 14,000 = $18,500 [Answer]

b.

Marginal tax rate is the rate of tax on the last dollar of income, which is 35% here.

c.

Average tax rate = (Total tax / Taxable income) × 100

                            = ($18,500 / $70,000) × 100

                          = 26.42857… %

                            = 26% (rounded)